OpenStudy (anonymous):

how can i know that any rational number that cannot be represented as terminating decimals?

OpenStudy (anonymous):

when you divide a number,if certain digit gets repeated again n again,then its repeating decimal,which can not be terminated...like 1/3 = 0.333333333333333333 and 3 keeps on repeating it can not be terminated got it?

OpenStudy (anonymous):

so -4/25 is terminating or not

OpenStudy (anonymous):

-4/25=-0.16

OpenStudy (anonymous):

it isnt at 6 it ends,you dont need to divide it anymore because the remainder is 0 when you dont get remainder 0 then its non terminating

OpenStudy (anonymous):

3/7=0.428571428 so what can we say about it

OpenStudy (anonymous):

still non terminating,as 428571 is getting repeated again?

OpenStudy (anonymous):

so if we are getting 0.636363 then its non terminating?

OpenStudy (anonymous):

yes if u r nt getting remainder 0 then its non terminating certain series of numbers or just a numner gets repeated

OpenStudy (anonymous):

number*